How many numbers of three digits can be formed with the digits 1, 2, 3, 4 and 5



if the digits in the same number are not repeated? How many such numbers are



possible between 100 and 10,000

1
Expert's answer
2021-12-26T17:47:28-0500

Solution:

The numbers of three digits can be formed with the digits 1, 2, 3, 4 and 5 if the digits in the same number are not repeated "=5\times 4\times3=60"


Three digit numbers: 5 "\times" 4 "\times" 3 = 60

Four digit numbers: 5 "\times"4 "\times" 3 "\times" 2 = 120

Five digit numbers: 0 (Because they’d be bigger than 10,000).

Total: 180 numbers